OpenStudy (anonymous):

Find (dy/dx) of √(xy)=x+3y help please!

OpenStudy (anonymous):

start with \[\frac{1}{2\sqrt{xy}}\times \left(y+xy'\right)=1+3y'\] and solve for \(y'\)

OpenStudy (anonymous):

Why would you start with that? I don't understand how you changed it to that.

OpenStudy (anonymous):

you are looking for the derivative with respect to \(x\) right? i.e. \(\frac{dy}{dx}\)

OpenStudy (anonymous):

That is correct.

OpenStudy (anonymous):

and the initial equation is, if i read it correctly \[\sqrt{xy}=x+3y\]

OpenStudy (anonymous):

That is also correct.

OpenStudy (anonymous):

ok then we are thinking that \(y\) is a function of \(x\) even though we don't know a formula for it, i.e. \(y=f(x)\) that is why it is called "implicit", because there is no explicit formula for \(f(x)\) so lets change all the \(y\)'s to \(f(x)\)

OpenStudy (anonymous):

Alright, I'm following...

OpenStudy (anonymous):

\[\sqrt{xy}=x+3y\] \[\sqrt{xf(x)}=x+3f(x)\]

OpenStudy (anonymous):

the derivative of the right side is easy, it is \[1+3f'(x)\] which i wrote more succinctly as \[1+3y'\] so far so good?

OpenStudy (anonymous):

Ok yes, I understand that now, Thanks. Continue,.

OpenStudy (anonymous):

the left side is the one that is a pain for that you need the product rule and the chain rule before we do it, suppose you had to take the derivative of \[\sqrt{x\sin(x)}\] would that be okay?

OpenStudy (anonymous):

Yes it would!

OpenStudy (anonymous):

ok you would get \[\frac{1}{2\sqrt{x\sin(x)}}\times (\sin(x)+x\cos(x))\] by the chain rule and the product rule

OpenStudy (anonymous):

okk

OpenStudy (anonymous):

similarly for \[\sqrt{xf(x)}\] you get the derivative as \[\frac{1}{2\sqrt{xf(x)}}\times (f(x)+xf'(x))\]

OpenStudy (anonymous):

replacing \(f(x)\) by \(y\) and \(f'(x)\) by \(y'\) you get \[\frac{1}{2\sqrt{xy}}\times (y+xy')\]

OpenStudy (anonymous):

Ohhh ok

OpenStudy (anonymous):

and then, like you said, solve for y'?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the algebra is going to be a drag, but it is algebra

OpenStudy (anonymous):

Thats not a problem! I'm pretty quick with Algebra.

OpenStudy (anonymous):

have fun hope it is more or less clear